If a hexagon has a radius (center to point of angle) of 6, what is the area of the hexagon?

Answer by rfer(10417)

The area of a hexagon is 6 times the area of each section.
the angle of each section is 60 degrees
radius = 6
sin60(6)=5.2 opposite side
a=(bh)/2
a=(5.2*6)/2
a=15.6 each section
a=15.6*6=93.6 sq units